Fermi National Laboratory
 

Fermilab Shutdown Report for August 28, 2009 - September 4, 2009

 

Fermilab Shutdown Report

8/28/09 – 9/4/09

LINAC

Shutdown Week 12 – On Schedule

  • Preaccelerator
    • Running on H- Source at 55.5 mA
    • Checked H- Ion Source brushes
    • Replaced TRIM90 power supply
  • I- Source Status:
    • The column was pulled apart
      • Plan:
        • Machine shop will polish up the electrodes.
        • Once the electrodes are back, buff them out, wash in sonic cleaner and boil them
        • Pre Acc experts will grind/polish the ceramic once again.
          • Estimate: 3 weeks
  • Low Energy
    • 750 KeV Line Buncher gradient bouncing problem
      • Problem returned last Sunday.
        • Replaced attenuators and mixer
        • Add new points to monitor the system
        • No bouncing was observed since then.
    • Replaced ion pump for LRF2 (HE side)
    • Linac RF station (LRF2) excessive crowbar period experienced last Monday evening. Forced conditioning of 1 out 3 Switch tubes. Currently running at lower voltage to change bias point
  • High Energy
    • Adjusted a Klystron RF station (KRF3) to improve the recovery time after a trip
    • Water group installed last set of heat exchanger plates on KRF1
    • KRF5 controls problem
      • Symptom: not inhibiting beam when station was off
      • Resolution: bad relay on the node
    • KRF3 body water flow meter replaced
  • Next Week
    • Support the Neutron Therapy Facility (NTF) patient treatment (Tuesday, Wednesday, & Friday)
    • Address problems as they reveal.
    • Provide beam to downstream machines.
    • Continue machine tuning
    • Investigate problems with crowbar recovery local application

Booster

Shutdown Week 12 – On Schedule

  • Activities
    • Running 5-6T ~2.5E12, 87% efficiency to MB
    • Tuning on 8T ~3.5E12, 85% efficiency
    • Accessed tunnel Thursday for small jobs during the MW810 work for MI
    • Possible drifting of MH2 400 MeV power supply, investigating
    • Continue tuning, reducing losses, improving efficiencies, and increasing intensity

Antiproton Source

Rings/Transport

  • Running 5-6T ~2.5E12, 87% efficiency to MB
  • Tuning on 8T ~3.5E12, 85% efficiency
  • Accessed tunnel Thursday for small jobs during the MW810 work for MI
  • Possible drifting of MH2 400 MeV power supply, investigating
  • Continue tuning, reducing losses, improving efficiencies, and increasing intensity

Beam Lines

  • Periods of controlled access only
  • Where able to turn on AP1 supplies (except trims)
  • Had an LCW flow issue in transport
  • Completed all critical device testing

Access

  • AP2 Beam Stop Repair
  • Pbar Rings access on Tuesday to install Anti-hydrogen beam line resulted in a vacuum being let up in three Accumulator vacuum sectors

Beam Stop (D:BSC700) Repair

  • Beam Stop was stuck closed.
  • Investigation revealed that the motor was turning by the beam stop was not moving
  • Nylon gear was missing a few teeth

Anti-Hydrogen Line Work

  • BV862 was installed over the shutdown between the Accumulator and the revived anti-hydrogen line
    • This used to be E862 back in 1996/97
  • After the faceplate was loosened to attach the anti-hydrogen beam line, vacuum was accidentally let up to air in three vacuum sectors.
  • Primary focus at this point was recovering the vacuum

Vacuum Recovery

  • Ran vacuum pumps overnight Tuesday/Wednesday, turned on ion pumps Wednesday morning, and launched straight into bake prep
  • BV862 was cycled and leak checked successfully
  • The bake is a five-day (plus setup) process
  • Originally thought we could not bake all three houses at once, but we were able to work around this.
  • The bake is in progress for A10, A50 and A60

Pbar Start Up

  • Friday – Sunday
    • Monitor the A10, A50 and A60 bakes
    • Fix any problems with the bake
    • Monday
      • Terminate the bake for A10, A50 and A60
      • De-gas vacuum components
      • Let tunnel devices thermally cool
      • Turn on ion pumps and turn off turbo pumps
    • Tuesday Morning
      • Reconnect all electronics in A10, A50 and A60
      • Reconnect liquid nitrogen to the A60 tanks and start cool-down
      • Final tunnel preparations before start-up
      • Troubleshoot known problems (such as hardware controls issues with A:BV607)
    • Tuesday Afternoon
      • Power Supply turn-on
      • Run usual turn on from access aggregate
      • Fix any problems encountered
    • Tuesday Evening
      • Ready for 8 GeV beam
    • When 8 GeV Protons are Available from MI
      • Reverse Protons to the Accumulator
        • Accumulator orbits, tunes, admittance
        • Debuncher orbits, tunes, admittance
    • When 120 GeV Protons are Available from MI
      • Target Station check-out
      • Beam to the Debuncher
        • RF system check-out
        • First pass at Debuncher cooling phasing (when cold)
      • Beam to Accumulator, stack pbars
        • Phase 4-8 GHz Core systems
      • Transfer pbars to Recycler
    • When Stable Operations Resume
      • Full check-out of all cooling systems

Main Injector

During the week we have assisted with establishing beam down the MI-8 line to MiniBooNE.

Tunnel Work

MI-40

  • Last Friday: Investigated a leak on the MI-40 Lambertson
    • Found small leak on the 18” downstream flange
    • Found smaller leak from corrosion on the beam pipe
    • Found smallest leak on the titanium window to the abort line
  • All leaks repaired and leak checked Tuesday after lunch and started to bake the Lambertsons
  • Bake is expected to run through Monday, requiring daily accesses to monitor
  • Four hour access expected on Tuesday to pull bake equipment
  • During MI-40 repairs we replaced two 600L ion pumps in the RF section
  • Finished up inspections and housekeeping
  • Thursday during the access to monitor the bake we inspected the 810 multiwire and 811 BPM. The BPM was fixed but the multiwire will require breaking vacuum to fix

Recycler

The Recycler is ready and waiting for beam.

Tevatron

Other Items

  • Luminosity counter at B0 replaced
  • All sign off done
  • Cold cathode read out problem holding off going to 4 K
  • 8 houses warmed to room temp
  • Leaks fixed in 9 houses
  • 88 magnets unrolled or moved
  • New vertical crystal collimator installed
  • Maintenance on many items in TeV

NuMI

Shutdown Week 12

  • The target hall work is done
  • Power supply testing over the weekend
  • Need to work on heat exchanger

Cryogenic system

TeV at LN2 Temperatures

Hydrocarbon Contamination Plugging in Heat Exchangers

Cryo Operations and Maintenance Status

  • Continued support AD/MSD for vacuum repairs: D3, E3, F1, A3, A2, A1, and Pbar Accumulator tanks are back at LN2 temperature. D1 and E1 are cooling down to LN2 temperature after repairs by MSD.
  • Contamination plugging of heat exchangers occurred last week in the LN2 precooled HTX-1/2. Contamination source is suspected as oil volatiles due to failed heater thermostat in A0#1 Mycom compressor; it has been isolated. Recent helium analysis indicated over 6000 ppm THC in this compressor. Eleven TEV FRIG heat exchangers were derimed to remove trapped contamination.
  • Pbar helium system cooldown to LHe temperatures was started but halted due to excessive heat exchanger pressure drop; exchanger derimed and cooldown initiated once again.
  • Pbar Accumulator LN2 tanks isolated for room temperature warm-up upon request.
  • Ongoing numerous FRIG/Compressor controls/instrumentation maintenance/repairs continue.
  • Ready for cooldown to LHe temperature upon approval

---------------------------------------------------------------------------------------

Tevatron cryo schedule is dependent on when we start system cooldown and problems encountered as such.

CHL HELIUM and NITROGEN PLANTS

  • 99.99% of all scheduled routine mechanical and electrical maintenance is complete

Current operations:

  • CHL Helium Cold box-2 operating with compressors “B” and “D”, online to LHe Dewars with LHe Pump-2.
  • Bulk LHe (13,000 pounds or 12,000 gallons or 50,000 liters) was delivered and ready for Tevatron FRIG system cooldown.
  • Mobile Helium Purifier at CHL has been successful to purify the Tevatron FRIG system, including the hydrocarbons

MiniBooNE

  • Taking Data
  • On 9/3/09 the vacuum pumps tripped off for unknown reasons. Everything turned on okay

Collider Detector at Fermilab

  • The Detector is closed up

D-Zero

EE Support

  • Replaced the NH2 power supplies
  • Closed two MI transformers
  • Will conduct final walk throughs today

FESS

Chillers

  • 55LCW Chiller #3 – tube & header work completed. Leak checked & eddy current tested. Re-assembling chiller work in progress. Under vacuum over weekend & will recharge with refrigerant next week. Tentative completion by Fri 9/11.
  • Comfort chiller #5 – refrigerant processing & charging completed. Expect to test run today & ready as back-up.
  • Chilled water operation – re-configured night set back controls to maintain chiller temperatures 24/7 for AD areas on a consistent basis until chiller #3 is on line.
  • Water treatment – adding additional new chemical treatment injection point for improved calcium carbonate control for condensers & heat exchangers.
  • Process chiller #2 – looking at scheduling this ~Dec to inspect & clean condensers

ICW

  • Zebra Mussels Spot flushing is completed.
  • Installing improved brominating sparge ring at the Casey’s pond pump inlet

Mechanical

  • The MI bake is going well

Instrumentation

Safety

Continue using the Take Five principles before, during, and after work activities to:

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/04/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory